OpenStudy (anonymous):

Suppose that GCD(a,b) = 1 and that a divides bc. Show that a must divide c without using the Fundamental Theorem of Arithmetic.

OpenStudy (anonymous):

I have questions about all of the proofs that I've found. For example, this one: Suppose a|c and b|c. Then c = am and c = bn. Suppose that gcd(a,b) = 1. We may write it as a linear combination: 1 = ax + by Observe that 1 = ax + by c = acx + bcy c = a(bn)x + b(am)y c = ab(nx) + ab(my) c = ab(nx + my). Hence, ab|c. Why does c = am and c = bn?

OpenStudy (anonymous):

Never mind, stupid question.

OpenStudy (rational):

its not a stupid question

OpenStudy (anonymous):

I think I get it.

OpenStudy (rational):

a|c and b|c a|c means `c = am` for some m b|c mean `c = bn` for some n

OpenStudy (rational):

example : 3|12 means `12 = 3*4`, here m = 4 2|12 means `12 = 2*6`, here n = 6

OpenStudy (anonymous):

That means that there is some integers for a and b which, when multiplied by a and b, will cause them to meet at a point c on the number line. That seems to be a given, but I'm trying to understand why.

OpenStudy (rational):

thats possible only when gcd(a, b) = 1

OpenStudy (rational):

for example : a = 4, b = 6, c = 12 clearly 4|12 and 6|12 but 4*6 does not divide 12

OpenStudy (anonymous):

Euclid's Lemma says that if a prime divides the product of two numbers, then it must divide at least one of those numbers. Going back to my original problem, nowhere do we know that either of the numbers are prime. We know that they're coprime, but one (or both?) of them could still be composite. Unless I'm missing something, Euclid's Lemma doesn't apply.

OpenStudy (rational):

Your original problem as-it-is itself is called Euclid's Lemma See http://prntscr.com/4j5i6u

OpenStudy (anonymous):

On the last line of that proof, how can a | (acx + bcy) be recast as a|c?

OpenStudy (anonymous):

Wait, never mind. That really was a stupid question, lol.

OpenStudy (anonymous):

Thanks, I understand!